hhanan:

for question 4, what you have is correct, you just need to replace the question marks with the correct number

hhanan:

hhanan:

@jhonyy9 could you help me?

jhonyy9:

what mean this ? from this radical you need getting 4 ? and for this you need write there the index of radical or what ?

jhonyy9:

this 4 in front of radical mean question nr. 4 ?

jhonyy9:

@TETSXPREME any idea please about this ?

jhonyy9:

@supie

jhonyy9:

do you understand this posted problem ?

TETSXPREME:

tbh idek is there 2 answers?

jhonyy9:

2 answers ? what is the question ?

hhanan:

25 7/2 with the original questions but then I answer \[\sqrt{25}^7\]

jhonyy9:

ok and what you need to do now ? do you need simplifie this radical or calcule this radical or what ?

jhonyy9:

so you can write again the 25 like square of how many ?

hhanan:

he just said that i need to replace to question marks with a number

jhonyy9:

ok but what mean this 4 in front of radical ?

jhonyy9:

the question wan be 25^(7/2) = ?

hhanan:

yes

hhanan:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 ok but what mean this 4 in front of radical ? \(\color{#0cbb34}{\text{End of Quote}}\) the 4 was just the number of the problem

jhonyy9:

ok than this is very very easy just rewrite the 25 like a power of 5

jhonyy9:

25 = 5^?

hhanan:

5^2

jhonyy9:

exactly

jhonyy9:

so you have there in this way (5^2)^(7/2) = ? what you need to do with exponents in this case ?

hhanan:

78125

jhonyy9:

\[(5^2){^{\frac{ 7 }{ 2 }}} = ?\] look please

hhanan:

5^7

jhonyy9:

perfect

hhanan:

so the 5 would have to go one top

jhonyy9:

no this mean 5 on power of 7

hhanan:

he said that the only thing that was missing was a number where the question mark is

hhanan:

hhanan:

(number 4)

jhonyy9:

hence is more understandably \[25^{\frac{ 7 }{ 2}} = ?\] in this case just you need to know that when there is a fractional exponent so the denominator of this fraction mean always the index of radical for example \[x^{\frac{ 2 }{ 3}} = \sqrt[3]{x^{2}}\]

jhonyy9:

ok ?

jhonyy9:

questions ?

hhanan:

so \[\sqrt[5]{25}^7\]

jhonyy9:

how you get this index of 5 ? there is on exponent 7/2 so 2 will be the index of radical

jhonyy9:

ok ?

hhanan:

\[\sqrt[2]{25}^7\]

jhonyy9:

\[\sqrt{25^{7}}\] in this way

hhanan:

but he said that there need to be a number

jhonyy9:

no the text of your posted problem said that write these exponential expressions in radical form

hhanan:

oh ok thank you

jhonyy9:

anytime yw